OpenStudy (sugarmochi):

Please help :c Find the coordiaees of the vertices of a triangle, formed by the linear graphs of y=-x+1, 2x-y=2 and 2y=x+2

OpenStudy (phi):

you should first try to find where the first two lines meet: y=-x+1 2x-y=2 one way is replace the y in the second equation with (-x+1), like this: 2x - (-x+1)= 2 can you "solve for x" ?

OpenStudy (sugarmochi):

would x = 1/4 ?

OpenStudy (phi):

can you show your work? 2x - (-x+1)= 2 becomes 2x +x -1 = 2 now what ?